OpenStudy (anonymous):

If (x1, y1) and (x2, y2) satisfy xy = k, then x1y1 = x2y2. Find x or y as indicated. (4.5, y) and (12, 6) a.y = -2 c.y = 16 b. y = 0 d. y = 12

OpenStudy (neer2890):

they satisfy x1y1=x2y2 so 4.5*y=12*6 y=72/4.5=16